weight of fuel .98g carbon 90 and hydrogen 8 what is gross calorific vlue and net calorific value whentemprature rise is 1.8 aadnwater eq 450 on assignment.com


1
Expert's answer
2021-11-10T00:30:14-0500

98-90= 8

8×8×1.8= 115.2

450/115.2= 3.91